A simple series circuit has an inductor of 1henry , a capacitor of 10^-6 farads and a resistor of 1000ohms . The initial charge on the capacitor is zero. If a 12volt battery is connected to the circuit and the circuit is closed at t=0, find the charge on the capacitor 1sevond later and the steady state charge.
1
Expert's answer
2020-03-16T10:56:31-0400

The voltage and the current relationships in circuit analysis


"i={dq \over dt}"

voltage drop across the inductor


"V_L=L{di \over dt}=L{d^2q\over dt^2}"

voltage drop across the resistor


"V_R=Ri=R{dq \over dt}"

voltage drop across the capacitor


"V_C={q \over C}"

Kirchhoff's Second Law


"L{d^2q\over dt^2}+R{dq \over dt}+{q \over C}=E(t)"

Given that "L=1\ H,\ R=1000\ ohm, \ C=10^{-6}\ F,\ E(t)=12\ V"


"1{d^2q\over dt^2}+1000{dq \over dt}+{1 \over 10^{-6}}q=12"

"{d^2q\over dt^2}+1000{dq \over dt}+1000000q=12"

The characteristic equation is


"r^2+1000r+1000000=0"

"D=(1000)^2-4(1000000)=-3(1000000)"

"r={-1000\pm i1000\sqrt{3} \over 2}"

"q_c=e^{-500t}(c_1\cos{(500 \sqrt{3}t)}+c_2\sin{(500 \sqrt{3}t)})"

Find the particular solution of the non-homogeneous differential equation using the method of the undetermined coefficients.

Assume that "Q(t)=A" is a solution of the non-homogeneous differential equation where "E(t)=12."

Differentiate the assumption with respect to "t"


"Q'=0,\ Q''=0"

Substitute in the original equation


"0+0+1000000A=12"

"A=1.2\times10^{-5}"

"Q(t)=1.2\times10^{-5}"

"q(t)=e^{-500t}(c_1\cos{(500 \sqrt{3}t)}+c_2\sin{(500 \sqrt{3}t)})+1.2\times10^{-5}"

The initial charge of the capacitor is zero


"q(0)=e^{-500(0)}(c_1\cos{(500 \sqrt{3}(0))}+c_2\sin{(500 \sqrt{3}(0))})+1.2\times10^{-5}=0"

"c_1=-1.2\times10^{-5}"

The circuit is closed at "t=0: i(0)=q'(0)=0"


"q'(t)=-500e^{-500t}(-1.2\times10^{-5})\cos{(500 \sqrt{3}t)}-""-500e^{-500t}c_2\sin{(500 \sqrt{3}t)}+""+1.2\times10^{-5}(500 \sqrt{3})e^{-500t}\sin{(500 \sqrt{3}t)}+""+c_2500 \sqrt{3}e^{-500t}\cos{(500 \sqrt{3}t)}"

"q'(0)=-500(-1.2\times10^{-5})-0+0+c_2500 \sqrt{3}=0"

"c_2=-{\sqrt{3} \over 1.2}\times10^5"

"q(t)=e^{-500t}\big(-1.2\times10^{-5}\cos{(500 \sqrt{3}t)}-{\sqrt{3} \over 1.2}\times10^5 \sin{(500 \sqrt{3}t)}\big)+""+1.2\times10^{-5}"

"Coulomb: [C]"


"q(1)=e^{-500(1)}\big(-1.2\times10^{-5}\cos{(500 \sqrt{3}(1))}-{\sqrt{3} \over 1.2}\times10^5 \sin{(500 \sqrt{3}(1))}\big)+""+1.2\times10^{-5}\approx1.2\times10^{-5}(C)"


Find the steady state charge.

"t\to\infin: q_p(t)\to1.2\times10^{-5} \ C"